OpenStudy (anonymous):

I am stuck again. The population of a city increases by 4,000 people each year. In 2025 the population is projected to be 450,000 people. What is an equation that gives the city's population p (in thousands of people) x years after 2010? I think it is p - 450 = 4(x - 15)???

OpenStudy (xapproachesinfinity):

year zero counts from 2010 find the slope of the population between 10 to 25

OpenStudy (xapproachesinfinity):

let's imagine that the population started from 10000! each year it will increased by 4000 so the next year we will have 14000, and so on to find the the increase of population per time which is the slope

OpenStudy (xapproachesinfinity):

You know how to get the slope don't you

OpenStudy (anonymous):

But I need to write an equation that gives the city's population. I came up with p - 450 = 4(x - 15) But I am not 100% on this

OpenStudy (xapproachesinfinity):

well that's what we are trying to come up with by finding the slope we can get the equation

OpenStudy (anonymous):

Yes I know how to find slopes but I'm not very good at it

OpenStudy (xapproachesinfinity):

m=450-4/25-10=?

OpenStudy (campbell_st):

you know the population isP = 450000, t= 15 years after 2010. You also know the growth rate 4000 = m so the equation is P = mt + i i is the initial population for the model. substitute the information and find i. then you have the model...

OpenStudy (xapproachesinfinity):

Oh forgot what i said the slope is 4000 like @campbell_st said

OpenStudy (xapproachesinfinity):

i was use the info incorrectly lol so you have m=4000 you only need to find the initial population

OpenStudy (anonymous):

60000

OpenStudy (anonymous):

the starting population would be 390000

OpenStudy (xapproachesinfinity):

how did you find it!

OpenStudy (xapproachesinfinity):

p=4t+i are you using this equation

OpenStudy (anonymous):

yes I am using that equation

OpenStudy (xapproachesinfinity):

so for t=25 the population is what

OpenStudy (xapproachesinfinity):

t=15 i mean

OpenStudy (anonymous):

I took 4000 and multiplied it by 15 that gave me 60000 then I took 450000 and subtracted 60000 to come up with 390000 for the starting population

OpenStudy (xapproachesinfinity):

hmm that would correct then!

OpenStudy (xapproachesinfinity):

so the equation is what?

OpenStudy (anonymous):

p = 4x + 15

OpenStudy (xapproachesinfinity):

how come? you just said the initial pop is 390000

OpenStudy (anonymous):

oh...sorry

OpenStudy (xapproachesinfinity):

2010 is the t=0 okay

OpenStudy (anonymous):

I hate this it confuses me sooooo much

OpenStudy (anonymous):

why is 2010 the t?

OpenStudy (anonymous):

I used 15 as the t to come up w/the initial population

OpenStudy (xapproachesinfinity):

y=4x+390 where y is the population in thousands and x is the number of years

OpenStudy (anonymous):

x is 15 because its after 2010 right

OpenStudy (xapproachesinfinity):

15 is you x no your y in the equation it the the number of years to get 450 thousands pop

OpenStudy (xapproachesinfinity):

correct

OpenStudy (anonymous):

is y the beginning population?

OpenStudy (xapproachesinfinity):

if i set x=0 which corresponds to 2010 i get 390 thousands which is the initial pop in the city in that year if i keep adding 4 after eah year i get my graph

OpenStudy (xapproachesinfinity):

No y represent the population in the city in and how it increases with time

OpenStudy (anonymous):

so y = 4000

OpenStudy (xapproachesinfinity):

No y changes with time it start from 390000 and it goes up

OpenStudy (anonymous):

in the equation p =mt + i then m=4000 t=15 and i =?

OpenStudy (xapproachesinfinity):

4000 is our change how fast is the pop changes with time they told it is changing by 4000 each year

OpenStudy (xapproachesinfinity):

i=390 in thousands t=15 corresponds to the year 2025

OpenStudy (anonymous):

I feel bad because you are helping me sooo much and I flat out stink at math

OpenStudy (xapproachesinfinity):

Eh don't worry mate^_^ little by little you get this

OpenStudy (anonymous):

what represents the m then?

OpenStudy (xapproachesinfinity):

it the how fast is our pop grows with time which they told is 4000 each year we represent it as 4 only in equations but we know it is 4 thousand if we graph is we would indicate that

OpenStudy (xapproachesinfinity):

it* sorry for grammar lol

OpenStudy (anonymous):

then I take m (4) multiply that by t(15) and then add i(390)

OpenStudy (anonymous):

It's okay about the grammar I can understand you! lol :)

OpenStudy (xapproachesinfinity):

why would you do that? it would give 450 in thousands which is the population in the year 2025

OpenStudy (anonymous):

Shoot! I thought I was looking for the population in years ??? I'm So Stupid!

OpenStudy (xapproachesinfinity):

you are looking for an equation that models the situation given to you^_^

OpenStudy (xapproachesinfinity):

and never feel that way! from mistakes we learn! just talking about this I'm learning too^_^

OpenStudy (anonymous):

well here are my choices..... p = 4x + 50 p - 450 = 4(x-15) p - 15 = 4(x -450) p = 4x + 15

OpenStudy (anonymous):

I thought is was the p - 450 = 4(x -15)

OpenStudy (xapproachesinfinity):

i see you are looking at the choices which confused you

OpenStudy (anonymous):

Thanks for being so kind!!! :)

OpenStudy (xapproachesinfinity):

yes it should be that one because it is the same as y=4x+390 lol

OpenStudy (anonymous):

Yes because when you were explaining and I was using the model P=mt + i I couldn't understand how any of the other 3 would work

OpenStudy (anonymous):

So I was right?!!!?? WOW

OpenStudy (xapproachesinfinity):

we did it differently than they did but the concept is what matters yes you were correct

OpenStudy (anonymous):

I didn't want to give the answer choices when I started this conversation because I don't want someone to just tell me the correct answer. I want to do my best to try to understand even if it is a little bit.

OpenStudy (xapproachesinfinity):

P=mt+i is the same as y=mx+b we are just changing names the equation you were given is y-y1=m(x-x1) know as point slope form

OpenStudy (anonymous):

I at least learned how to come up with the initial population which was helpful to me. Thank you again for you patients and for helping! Happy Thanksgiving!! :)

OpenStudy (xapproachesinfinity):

the principal is the key! once you know it you can play with it however you like

OpenStudy (anonymous):

I will work with it. and see how I do. :) I'm afraid math is not going to be in my career future!! :D

OpenStudy (xapproachesinfinity):

np! and thanks happy thanksgiving to you as well^_^ keep up with the good work

OpenStudy (xapproachesinfinity):

math is actually plenty of fan! you just need to be open minded about it lol

OpenStudy (anonymous):

You will make a good teacher someday! :)

OpenStudy (xapproachesinfinity):

^_^ Not as good as some wonderful minds here in OP ============ here is a similar problem to this i started with 1000$ in an account, my money is decreased by 50$ a year in what year i would owe 2000$

OpenStudy (xapproachesinfinity):

work on this for later^_^ you will see that you can in fact do this belivee me

OpenStudy (xapproachesinfinity):

oh I forgot to add i have 1000$ in the account in 2007

OpenStudy (xapproachesinfinity):

there you have the starting time^_^

OpenStudy (anonymous):

it will be in 2047

OpenStudy (anonymous):

I think you need to put your money in a investment account so you can gain 50 a year instead of loosing it! :D lol

OpenStudy (anonymous):

Don't ask me to write a formula because that will take us another hour!!! :o)~

OpenStudy (xapproachesinfinity):

hmm i just came up with the problem for fan ^_^ surely i wouldn't like to loose money lol

OpenStudy (xapproachesinfinity):

No need! i have to go nice talking to you^_^

OpenStudy (anonymous):

lol .... okay! Thanks Again! Take Care! :o)

OpenStudy (xapproachesinfinity):

Take care^_^